OpenStudy (anonymous):

Can Some Check My Answer? Which of the following is the slope between the two points (-1, -1) and (8, 0)? A. 9 B. -9 C.1/9 D.-1/9 I Believe That the answer might be a because the 8 its a negative ?

jimthompson5910 (jim_thompson5910):

did you use the slope formula m = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

isnt*

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) m = (0 - (-1))/(8 - (-1)) m = (0+1)/(8+1) m = ??

OpenStudy (anonymous):

i just did it and i got -1/9

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

nvm its 9/1 no negative so the answer is C

jimthompson5910 (jim_thompson5910):

it's 1/9, so yes, C

OpenStudy (anonymous):

@jim_thompson5910 Thanks Bro For putting up the steps *really needed it*

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

could u check my new question please @jim_thompson5910